OpenStudy (anonymous):

EASY! PLEASE HELP. Find the equation of the line that is parallel to x=5y-9 and passes through (3,-6). Express answer in standard form.

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

@hba

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@TuringTest

OpenStudy (anonymous):

give me your number then ill help you

OpenStudy (anonymous):

lmao

OpenStudy (anonymous):

nice try

OpenStudy (anonymous):

your lost

OpenStudy (anonymous):

you're* ?

OpenStudy (anonymous):

lol.... I will help

OpenStudy (anonymous):

thanks ven

OpenStudy (anonymous):

The slop of the line it is parallel to is 5. Using slope intercept equation Y=5x+b Plug in the (x,y) given and solve for you intercept, can you do that?

OpenStudy (anonymous):

i already got y= -1/5x+33/5

OpenStudy (anonymous):

for the slope intercept

OpenStudy (anonymous):

This what i get -6=5(3)+b -6=15+b b=-21 So your intercept is -21

OpenStudy (anonymous):

I am not sure where you get -1/5 for your slope from, if it is parallel to the other one it have the same slope which is 5.

OpenStudy (anonymous):

NEVERMIND i messed up

OpenStudy (anonymous):

isn't that perpendicular

OpenStudy (anonymous):

lololol yeah

OpenStudy (anonymous):

You are right i was wrong i looked at it y= when it is x= let me fix it!

OpenStudy (anonymous):

alrighty!

OpenStudy (anonymous):

y=1/5x+9 so Y=1/5x+b -6=1/5(3)+b -6=6/10+b b=-6.6 Y=1/5x-6.6

OpenStudy (anonymous):

isn't it -1/5x?

OpenStudy (anonymous):

It would be -1/5x if it was perpindicular but you said the line is parallel to it

OpenStudy (anonymous):

okay true

OpenStudy (anonymous):

Sorry i mess up there before >.<

OpenStudy (anonymous):

where'd you get the 9 from?

OpenStudy (anonymous):

wait jk

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

The 9 is not important. It is the y-intercept of the equation it is parallel to

OpenStudy (anonymous):

i know i realized that lol

OpenStudy (anonymous):

You get it now? :P

OpenStudy (anonymous):

but also 1/5 multiplied by 3 is 3/5

OpenStudy (anonymous):

well so do i write that ending equation in standard form and thats it?

OpenStudy (anonymous):

Yeah 3/5=6/10, you had it right before me just made it perpindicular. Yeah that is the standard from

OpenStudy (anonymous):

no because standard form is ax+by=c

OpenStudy (anonymous):

me so confuzzled

OpenStudy (anonymous):

Oh no that is the standard form we alway use. sorrry i am making it worse. you can move them over to get -6.6=y-1/5x

OpenStudy (anonymous):

Which is -1/5x+y=-6.6 that is standard form so sorry

OpenStudy (anonymous):

oh yay! its right ! thanks bye u rock